OpenStudy (anonymous):

Which answer shows this equation in standard form? 8-4(x-y)--6x+6 A. -10x+4y=-2 B. 2x+4y=-2 C. 2x+4y=2 D. 2x-4y=-2 @superman38

OpenStudy (anonymous):

Your equation is messed up what --6

OpenStudy (anonymous):

oh it is 8-4(x-y)=(-6x)+6

OpenStudy (anonymous):

A linear equation formula is Ax+By=C Simplify the right hand side \[8-4(x-y)=-6x+6\] Simplify the left hand side \[8-4x+4y=-6x+6\] Add -4x and 6x to get 2x \[8+2x+4y=6\] Move all terms to the right hand side of the question that don't contain a variable. \[2x+4y=-8+6\] Then Add -8 and 6 to get -2. \[2x+4y=-2\] So C is your answer. @17Trinity_17

OpenStudy (anonymous):

not c sorry about that

OpenStudy (anonymous):

OMG B

OpenStudy (anonymous):

I dont know whats going on lol

OpenStudy (texaschic101):

lol....I was wondering about that for a sec